need to solve using graphing system
y=-3x+4
y=3x-2

Answer by mananth(16946)

y=-3x+4
when x= 0, y=4 (0,4)
when x= -1, y = 7 ( -1,7)
y=3x-2
when x=0, y=-2 (0,-2)
when y =1, x=1 (1,1)
Plot the points and draw the lines. The point of intersection is the solution.(1,1)



m.ananth@hotmail.ca